Is the function f (x) = |x^2 - x| differentiable at x = 2. If yes find it’s derivative.

Grade:upto college level

3 Answers

ruchi yadav
10 years ago
Yes fuction will be differentiable at x =2. which can easily be seen by graph of the function.

First draw the graph of y = x2-x and then take mode.

Derivative will be = 2x-1
at x= 2
Derivative will be 4-1 = 3


Thank You
Ruchi
Askiitians Faculty
Sher Mohammad IIT Delhi
10 years ago
yes,
for x>1 f(x)=x^2-x
f'(x)=2x-1
f'(2)=4-1

Sher Mohammad
B.Tech, IIT Delhi


bharat bajaj IIT Delhi
10 years ago
Yes, for x>1, the f(x) = x^2 - x
It if a polynomial function. Hence it is differentiable at x=2.
Thanks & Regards
Bharat Bajaj
askIITians Faculty